OpenStudy (blowery):

Help me solve x^2 + 2x -3 = 0

OpenStudy (hxkage):

Alright, do you know splitting the middle term method? It could be used here..

OpenStudy (blowery):

My first step was (X + 3) (X - 1) = 0; second step was (X + 3) = 0 and (X - 1) = 0; answer was X = -3 x = 1. Is that correct?

OpenStudy (hxkage):

Yup, you are right! Good work! :DD

OpenStudy (blowery):

I'm very new. How do I thank you?

OpenStudy (hxkage):

You could medal me by clicking on the medal icon.. U see that? :)

OpenStudy (blowery):

Do you mind me asking one more problem?

OpenStudy (hxkage):

Sure, ask me. :D

OpenStudy (blowery):

The area of this rectangle is 80ft^2. What is the length and width? One side is X. The other side is X + 16. I started with 80 = x (X + 16)

OpenStudy (hxkage):

That's right. :)

OpenStudy (hxkage):

What would be your next step? c:

OpenStudy (blowery):

If that's right, then I get 80 = x^2 + 16x

OpenStudy (hxkage):

I think you will have to take 80 to the other side and do what u did earlier.. Not sure.

OpenStudy (blowery):

Ok, I'll try that! Thank you

OpenStudy (hxkage):

No problemo, mate! :)

OpenStudy (blowery):

Length is 20, width is 4

OpenStudy (hxkage):

Right! :D

OpenStudy (hxkage):

Same method, ayee?